OpenStudy (anonymous):

cos(pi-o)+ sin(pi/2 +0)=0 verifying the identity

OpenStudy (jdoe0001):

http://frank.harvard.edu/~jason/APCalc/Image57.gif <--- expand the functions using the sum identities

OpenStudy (anonymous):

cos Pi = -1 ; and sin Pi/2 = 1, then cos (Pi -0) + sin (Pi/2 + 0) = 0

OpenStudy (anonymous):

cos(pi-o)+ sin(pi/2 +0)=0 LS= cos(pi-o)+ sin(pi/2 +0) LS=cos(pi)+sin(pi/2) LS=0+0 therefore 0=0, LS=RS

OpenStudy (anonymous):

because...:

OpenStudy (anonymous):